Magnesium metal reacts with hydrochloric acid to produce magnesium chloride in solution and hydrogen gas. How many hydrogen molecules are produced when 60.0 g of magnesium is reacted in excess hydrochloric acid? Show step-by-step solution.
1
Expert's answer
2020-05-26T14:21:49-0400

Solution:

The balanced chemical equation:

Mg(s) + 2HCl(aq) = MgCl2(aq) + H2(g)

According to the chemical equation: n(Mg) = n(H2)


Moles of Mg = n(Mg) = Mass of Mg / Molar mass of Mg

The molar mass of Mg is 24.305 g/mol

n(Mg) = (60.0 g / 24.305 g/mol) = 2.4686 mol


n(Mg) = n(H2) = 2.4686 mol

At STP, the volume of 1 mole of a gas contains 6.022×1023 atoms or molecules of that gas.

So,

Number of molecules of H2 = (2.4686 mol × 6.022×1023) = 1.4866×1024 = 1.49×1024


Answer: 1.49×1024 molecules of H2 are produced.